A motorcycle traveling 60mph overtakes a police patrol car traveling at 40mph that had a 3hr head start. How far from the starting point are the two vehicles?
a) 360miles b) 279 miles c) 187 miles d) 290 miles

Found 2 solutions by scott8148, uday1435:
Answer by scott8148(6628)

the car has a 3hr head start ___ at 40mph, the distance is 120mi (3*40)

the motorcycle catches up at a rate of 20mph (60-40)

it takes 6hr (120/20) for the motorcycle to catch the car

the motorcycle (and the car) are 360mi (60*6) from the starting point when the motorcycle catches up

Answer by uday1435(57)

Since the police car has a 3hr headstart, and is moving at 40 mph, he is 3 x 40 = 120 miles ahead of the motorcycle. To catch up this gap he has a relative speed of (60-40 = 20mph).
So the time needed to catch up = distance/speed = 120/20 = 6 hours.
In that time he has to travel a distance = speed x time = 60 x 6 = 360 miles
Amswer: 360 miles


If you still need further clarifications please contact udayakumar.t.r@gmail.com